Emily is mowing her lawn which is in the shape of a rectangle with a length of 10 feet and width of 12 feet. Emily completed mowing it in 3 hours.
What is the unit rate at which Emily mowed her lawn?

Answer by TimothyLamb(4379)

a = 10 * 12 = 120 sq.ft
---
r = a / t
---
r = 120 / 3
---
r = 40 sq.ft/hr
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php